Sol Summary – Nov 9th

[title Sol Summary – November 9th]

[category sol-summary]

Sol:13
Summary Title: New Crew Arrives
Author’s name: Guy Murphy
Mission Status: Simulation mode ended first light this morning.
Sol Activity Summary: Two of the crew members travelled to Grand Junction this morning to collect the new crew. After stopping at Hanksville for lunch, the four new crew members Steve Whitfield, Jennifer Lane, Shane Usher and Larissa Wilson arrived at the MDRS at around 3pm. The remainder of the afternoon was spent settling them into the campus. In the evening, the crew put on a special dinner for the new crew, Shannon, Atila and David.
Look Ahead Plan: Tomorrow Sandy Dance and Dianne McGrath will depart the station and orientation of the new crew will occur in preparation for commencing the full simulation on Monday morning.
Anomalies in work: –
Weather: Another sunny, clear day on Mars
Crew Physical Status: Crew in excellent health.
EVA: None undertaken
Reports to be filed: Sol Summary, Operations Report.
Support Requested: None

Operations Report – Nov 9th

[title Operations Report – November 9th]

[category operations-report]

Crew 214 Operations Report 9-11-2019
SOL: 13
Name of person filing report: Andrew Wheeler
Non-nominal systems: Opportunity, Suit 10
Notes on non-nominal systems: Opportunity brakes do not work, Suit 10 is not functioning
Generator: run
Hours run: 13h
From what time last night: 1815
To what time this morning: 0715
List any additional daytime hours when the generator was run: N/A
Solar— SOC % (Before generator is run at night): 70%
Diesel Reading – 55%
Station Propane Reading – 76%
Ethanol Free Gasoline: N/A
Water (loft tank): 41 gallons
Water Meter: 1456896 units
Water (static tank): 150 gallons (not filled)
Static to Loft Pump used – yes
Water in Green Hab: 225 gallons
Water in Science Dome: 0 gallons
Toilet tank emptied: yes
Deimos rover used: not at hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: no
Hours: 109.2
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging): 100%
Currently charging: yes
Opportunity rover used: no
Hours: 65.6
Beginning charge: 100%
Ending charge: 100%
Currently charging: no
Curiosity rover used: no
Hours: 115.0
Beginning charge: (Before EVA) : 100%
Ending charge: (On return from EVA, before charging): 100%
Currently charging: yes
Notes on rovers: no
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): no
Reason for use: N/A
Oil Added? no
ATV Fuel Used: 0 Gals
# Hours the ATVs were used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? N/A
CrewCar used and why, where? CrewCar to Grand Junction and return for Crew 215
General notes and comments: N/A
Summary of internet: Functional
Summary of suits and radios: Nominal
Summary of Hab operations: All systems nominal
Summary of GreenHab operations: Atmospheric equalization to the green hab was carried out during the day, plants were watered. The heater automatically came on at 5pm.
Summary of Science Dome operations: All systems nominal.
Summary of RAM operations: Not used
Summary of any observatory issues: Nominal
Summary of health and safety issues: All crew healthy
Questions, concerns and requests to Mission Support: Static tank not filled. Water heater unusable. Opportunity is currently unusable.

Operations Report – Nov 9th

[title Operations Report – November 9th]

[category operations-report]

MDRS Supplemental Operations Report 8-NOV-2019

Name of person filing report: Shannon Rupert

Reason for Report: Work done by staff for beginning of the season

Non-nominal systems: n/a

Action taken for non-nominal systems: n/a

Generator check, note if oil and coolant added: coolant added. We
still have a very small diesel leak to the generator and are
investigating a possible oil leak in the engine.

ScienceDome AC unit, on and operational: nothing to report

Solar— Running nominally with nighttime generator backup

Diesel Reading – % not checked

Propane Reading – not checked

Ethanol Free Gasoline – n/a gallons.

Water (director’s tank) – n/a gallons

Water (static tank) – n/a gallons

Water (GreenHab) – n/a gallons

Water (loft) – This tank was cleaned today. Hose was replaced and the
leak fixed. We still need to cushion the hose so it doesn’t develop a
kink like the last one

Static to Loft Pump used – yes

Water Meter: n/a

Toilet tank emptied: nothing to report

Deimos rover used: Still not operational

Hours:

Beginning charge:

Ending charge:

Currently charging:

Sojourner rover used: Yes, operation nominal

Hours:

Beginning charge:

Ending charge:

Currently charging:

Spirit rover used: Yes, operation nominal

Hours:

Beginning charge:

Ending charge:

Currently charging:

Opportunity rover used: Yes, brakes still an open issue. Buildup of
leaves from when it was in town caused an acid burn problem on the
batteries but they were cleaned today.

Hours:

Beginning charge:

Ending charge:

Currently charging:

Curiosity rover used: Yes, operation nominal.

Hours:

Beginning charge:

Ending charge:

Currently charging:

Notes and action taken on rovers: Names not replaced yet.

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3). No. 300 still in town.

Reason for use: n/a

Oil Added? n/a

ATV Fuel Used: 0 Gals

# Hours the ATVs were Used: none

Notes and action taken on ATVs: n/a

HabCar running? If used, why, where? HabCar was used to haul water
then taken for service on 5 Nov. Oil change and tire rotation were
completed but back brakes and wheel damage also needed repair, so it
is still off campus. To be retrieved on Monday.

CrewCar running, If used why, where? Yes, used to transport crew and
supplies back and forth to town and Grand Junction. It was off campus
from 5-8 Nov and had an oil change, tire rotation and check engine
light serviced. It was also cleaned inside and a full service
exterior wash and wax was done.

Campus wide inspection, if action taken, what and why? We removed a
bunch of ethernet cord running from the Hab to RAM. We suspect it was
part of a set up for the webcam but it was strung haphazardly and so
it will need to be replaced properly. Unfortunately it sustain some
damage and may need repair or replacement

Evidence of rodents, where and action taken: We seem to have more
rodents than we have had in years. Visitors are captured and removed
to the north. No house mice, only larger mice.

Summary of internet: All nominal. Business account needed a new
router/modem and was repaired on 4 Nov.

Summary of suits and radios: Nothing to report

Summary of Hab operations: New kitchen hardware arrived and was
delivered to Hab.

Loft water tank cleaned and repaired. The hose on top was replaced by
a new 4-foot RV hose. We tried to create a new electric plug for the
static tank, as the previous one melted last season and a temporary
cord has been used since then. We determined that the cord running
the pump was still safe and so we left it alone and will create a new
one for the tank heater running from the rear airlock plug. Not ideal
but it will work until the pump plug fails and then the system should
be replaced with an actual outlet. Someone plugged in the tank heater
while it was not in the water and so we need a new one.

We removed the two black office chairs installed by Ikea as part of
the refit and reinstalled the two sleeper loungers that had been
there. Crews use the loungers as a comfortable bonding space and the
two chairs were not conducive for that use. Covers have been ordered
to improve their outward appearance. We couldn’t put all four of them
back in place as the kitchen has been extended about 8 feet into that
space. We may put a third one back it we can find the hardware for
it.

Summary of Outpost operations: The intern trailer now has water- hot
and cold, heat, and wastewater tanks all operational. A propane tank
was installed at the front of the trailer, the same size as the one
for the director’s trailer. We fixed a small leak in the freshwater
tank and are finishing up with the cleaning/organization of the
trailer. We have not been able to get the refrigerator running on
propane. We have connected a 30 amp service from the power pedestal
to the trailer and everything seems nominal. The yard was enlarged so
the dogs would have more space.

Summary of GreenHab operations: Planting continues.

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Summary of health and safety issues: Nothing to report

List and summary of projects:

Tennis balls to protect the tarps on the tunnels: We purchased used
tennis balls to cap the t-posts in the tunnel instead of replacing the
urine cups we used to use as they fly away too easily with the wind.
It is very hard to cut the balls and even harder to get them on the
t-posts. But they should keep the tarps from ripping once the winds
start.

New food supply store at Rock Shop: We have received most of our food
supplies and have settled into our new and upgraded space. Some
things are still missing and/or the wrong size, but the funniest thing
was the amount of flour we received. Let’s just say we won’t run out
soon!

Questions, concerns, supplies needed and requests: Nothing at this time.